OpenStudy (anonymous):

Alan is twice as old as Sue and half as old as Joseph. If the average (arithmetic mean) of all three ages is 14, how old is Sue? 1. 6 2. 7 3. 14 4. 16 5. 21

OpenStudy (anonymous):

2x=y x=2z 4z=y x+2x+1/2x=14(3) 3x+1/2x=42 7/2x=42 7x=84 x=12 y=6 z=24

OpenStudy (anonymous):

sue is 6

OpenStudy (anonymous):

thx